OpenStudy (anonymous):

Find the angle the measure describes. The angles measure is 22 degrees more than it's complement

OpenStudy (photon336):

If I remember this complimentary angles sum up to 90 degrees. so let's call our first angle C and our second angle theta. we know that theta is 22 degrees more than C so we have theta = 22 + C \[\theta = 22+C \] Compliment = C \[22+2C = 90 \] -22 -22 = 68/2 = 34 Compliment = 34 34 + 22 = 56 = theta 56 + 34 = 90

OpenStudy (anonymous):

Thank you! I didnt know how to do it ! or set it up.